Quantifierfree Stability in Generic Structures

Ingo Kraus

I will present a framework for local stability in nonelementary classes of ,,generic structures”. This generalizes standard stability theory as well as some remarks of Hrushovski’s about quantifierfree stability in universal domains and it also fits to the approach of Pillay in his preprint ,,Forking in the category of existentially closed structures”. As an example one has the class of generic automorphisms of a stable theory as a quantifierfree stable (but unstable) class of structures.

Address: Institut für mathematische Logik, Eckerstr 1, 79104 Freiburg, Deutschland

E-mail: kraus@mathematik.uni-freiburg.de